y'' - 3y' - 10y = 0 ; y(0)=1 and y'(0)=10


1
Expert's answer
2022-06-07T08:21:36-0400

Corresponding (auxiliary) equation


"r^2-3r-10=0"

"r_1=-2, r_2=5"

The general solution of the homogeneous differential equation is


"y=c_1e^{-2x}+c_2e^{5x}"

"y(0)=c_1+c_2=1"

"y'=-2c_1e^{-2x}+5c_2e^{5x}"

"y'(0)=-2c_1+5c_2=10"

"c_1=-\dfrac{5}{7}, c_2=\dfrac{12}{7}"


"y=-\dfrac{5}{7}e^{-2x}+\dfrac{12}{7}e^{5x}"